OpenStudy (anonymous):

COuld someone help me on this question, Find the derivative: f(x) = 3x^5-5x^3+4

OpenStudy (anonymous):

f'(x) = 15x^4 -15x

OpenStudy (anonymous):

sorry, second term is -15x^2

OpenStudy (anonymous):

Any chance you could tell me the steps? thanks

OpenStudy (turingtest):

f(x)=ax^n f'(x)=nax^(n-1) so for f(x)=3x^5-5x^3+4(x^0) f'(x)=5(3x^4)-3(5x^2)+0=15x^4-15x^2

OpenStudy (anonymous):

thanks